A point moves so that its displacement s is given by s=16+48t-t^3, when and where will it stop and reverse its direction of motion?
Answer by Alan3354(69443)

A point moves so that its displacement s is given by s=16+48t-t^3, when and where will it stop and reverse its direction of motion?
------------
Assuming t >= 0
---
s=16+48t-t^3
s'(t) = 48 - 3t^2 = 0
t^2 = 16
t = 4